# Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data

### v2

@@ V-D Summary
In this section we demystify the representative approaches to knowledge filtering within a federation. With rising concerns of data-centric attacks, accurately identifying adversarial clients that train on poisoned local data can be a top priority with regard to system security and robustness. We argue that fine-grained filtering at sample level can be employed jointly with client selection to provide more reliable evidence for adversary detection. As of data re-balancing, existing approaches stand on an implicit assumption that uniform data distribution is ideal for local training. However, this may not always hold. On the one hand, the local learner could under-fit on the “special knowledge” in local data when more attention is paid to the tail classes [76]. On the other hand, there is no sufficient theoretical evidence that supports imbalanced sampling, provided that the test data distribution is unknown. Although extensive effort has been made in addressing class-imbalanced learning, collaborative knowledge filtering for regression tasks, e.g. drug response modeling, is hardly explored. Moreover, existing studies pay little attention to the subtle line between data imbalance and non-IID data, the understanding of which can help us better reshape the knowledge locally and globally. In Table III we summarize the strength and limits of existing studies in relevance to knowledge filtering in the federation.
VI Label and Feature Space Correction
In this section we review the advance of research on label and feature space correction, by which we refer to the branch of studies that employ collaborative mechanisms to make correction or transformation on data labels or features. Provided that the quality of local data can be dubious especially in a cross-device federated system, it is important to rectify the samples in each local domain. In contrast to knowledge filtering that aims to exclude misinformation, amendments of samples usually involves data engineering techniques and feature engineering methods. Hence, in the following content we survey the state-of-the-art under two different categories: federated data correction and federated feature engineering.
@@ VI-A Federated Data Correction
As discussed in Section V, data noise commonly exists in a federation. Among various types and forms of noise, mislabeled samples are believed to have salient impact on supervised learning. Many researchers endeavored to address the problem of noisy labels in the context of FL, that is, federated noisy label learning (FNLL) [84, 85, 86, 87]. Considering the tie between a client and its local domain, it is intuitive to first identify the “noisy clients” [88, 89]. Xu et al. [88] proposed FedCorr, a multi-stage label denoising framework that streamlines a pre-processing stage and a fine-tuning stage prior to the traditional FL process, all executed in a federated fashion. In the federated pre-processing stage, two distinct mechanisms are employed for detecting noisy clients and correcting noisy labels. On the one hand, each client computes an estimate of the Local Intrinsic Dimensionality (LID) based on its local model’s predictions, which are collected by the central server for noisy client detection. As an useful measure that connects dimensionality and discriminability [90, 91], the LID for a vector $\boldsymbol{x}\sim\mathcal{D}$ is defined as:
@@
where $d_{x}$ is the distance between $\boldsymbol{x}$ and a randomly drawn vector $\boldsymbol{x}^{\prime}$ from $\mathcal{D}$ and $F_{d_{x}}(\cdot)$ represents the cumulative distribution function (CDF) of $d_{x}$. As models trained on noisy data tend to report larger LID scores, the server then uses this rule to spot these clients for label correction.
FedCorr [88] adopts a confidence-based rule to find and correct mislabeled data on noisy clients. Specifically, the subset of samples (controlled by a proportion $\pi$) that yield the highest losses is first selected, following which the predictions of the global model $f_{G}$ are used for relabeling. Formally, the final subset of samples $D^{\prime}_{k}$ (for a noisy client $u_{k}$) to be relabeled is defined in (11):
$D^{\prime}_{k}=\big{\{}(\boldsymbol{x},y)|(\boldsymbol{x},y)\in\underset{\hat{D}\subseteq D_{k}\atop|\hat{D}|=\pi|D_{k}|}{\arg\max}\sum_{(\boldsymbol{x},y)\in\hat{D}}\ell(f_{G}(\boldsymbol{x}),y)\\ \text{ and }\max f_{G}(\boldsymbol{x})\geq\theta\big{\}},$ (11)
where $\ell$ is the loss function and $D_{k}$ is the local data from client $u_{k}$. The labels of $D^{\prime}_{k}$ are then corrected to the predictions by the global model $f_{G}$.
There are two major defects in FedCorr [88]: i) incompatibility to general loss functions other than cross entropy, and ii) susceptibility to dishonest clients who tamper with their LID scores. Targeting at the second issue, Wang et al. [89] built upon FedCorr and redesigned its federated pre-processing stage to achieve robust label correction. The key idea behind is to craft an efficient Zero-Knowledge Proof (ZKP) [92] protocol by which the server (as the verifier) can verify the computation integrity of the LID scores reported by the clients (as the provers) in a batched manner. The proofs are generated both before and after label correction. Only calculation in the form of arithmetic circuits is verifiable by the protocol, which limits its potential application to other data correction schemes. Similarly, the FedELC framework by Jiang et al. [84] detects clients that carry noisy data and corrects them in two separate stages. The first stage is almost the same as in [88, 85], whilst the second stage creates learnable soft labels for data correction optimized towards the minimization of a linear combination of three different entropy losses. FedELC was evaluated on a rich set of noise data settings including the CIFAR-N [93] dataset from Amazon Mechanical Turk which features a high noise rate of human-annotated labels.
Noisy label detection and correction can also be executed directly at sample level. Based on label uncertainty estimation [94], Zeng et al. [95] started from utilizing a well-trained global model’s predictions to estimate the probability that samples of a class $i\in\mathcal{C}$ are correctly labeled in a local dataset $D_{k}$:
@@
where $D_{k,y=i}$ is the data subset on client $u_{k}$ labeled as class $i$, $\mathcal{C}\subset\mathbb{W}$ is the label space, and $\theta_{G}$ denotes the global model’s parameter. Next, the global estimate of the probability of correctly-labeled class $i$ can be obtained using weighted average consensus [95] over the population $\mathcal{K}$:
@@
which resembles the model aggregation rule in the canonical FedAvg [96].
A practical concern of federated data correction is that the participants can be self-interested, which means that they may not stick to the protocol if their gain of participation cannot cover the cost of data correction. In light of this issue, Yan et al. [97] first modeled the payoff of clients as a function of the model’s accuracy, payment from the task publisher, and the data correction cost. It is worth noting that the authors employed an iterative denoising strategy that corrects a fraction of $\alpha$ noise samples per round. To reduce the fraction of noisy data from $\epsilon$ to $\epsilon^{\prime}$ given a denoising cost per sample $c$ and data size $|D_{k}|$, the total cost of data correction for client $u_{k}$ is:
@@
Yan et al. [97] considered two different situations of federated denoising. In the first case, clients are cooperative and work towards to the optimal social welfare, i.e., the maximization of the sum of client payoffs. In the second case, every client is assumed self-interested and only motivated to maximize its own payoff, which corresponds to a non-cooperative game.
@@ VI-B Federated Feature Engineering
Prior to the boom of deep learning, feature engineering had proved very useful in alleviating the “curse of dimensionality” and allowing the model to learn more generalizable knowledge from the data. In the context of KAF, by federated feature engineering we refer to a new thread of research with particular focus on collaborative knowledge augmentation by means of aligning, selecting and transforming the feature space (Fig. 6). This process can either be independent of the downstream task or somehow coupled with a specific federated learning protocol. In both cases, feature engineering across multiple domains in the federation poses new challenges in several aspects.
  • Disordered or misaligned attributes: in cross-institutional cooperation it is common to deal with different views of data where the local datasets are subject to different permutations of attributes [71].
@@
These challenges have catalyzed rising interest in federated heterogeneous learning and also motivated cutting-edge studies on new feature engineering paradigms. In this section we discuss the advance in federated feature engineering, with particular focuses on collaborative feature selection and feature transformation.
Figure 6: A schematic illustration of collaborative feature engineering under the KAF paradigm.
Fed-FiS [98] represents an early exploration of collaborative feature selection in the federation based on information theory. The study follows a training-free, task-agnostic approach to assessing the importance of features. Locally, a client $u_{k}$ independently evaluates the feature-feature mutual information (FFMI) and the feature-class mutual information (FCMI) over the local feature and label set $\{f^{(k)}_{1},f^{(k)}_{2},\ldots\}\cup C^{(k)}$ by Eqs. (15) and (16):
@@
Based on these mutual information-based measures, each client performs clustering to select a subset of features which are then uploaded to the central server. In the second stage, the server uses FFMI and FCMI again to score each feature, which in turn helps the clients identify the best local feature subset. Banerjee et al. [99] further extended this framework using an improved ranking scheme based on Pareto optimization where each feature is treated as a solution characterized by FFMI and FCMI. The features are then ranked based on the relation of domination and reported back to the clients.
@@
Fu et al. [100] studied the feature selection problem in the Vertical Federated Learning (VFL) setting where the set of entities are associated with consistent identifies but represented by different features. The proposed framework, FEAST, involves three types of clients for collaborative feature selection: a selected party, an active party, and candidate parties. In the pre-processing stage, discretization is first applied to each feature, combined with stratified sampling to reduce computation and communication costs. This is followed by iterative operations and interactions between the parties. Specifically, a party is selected to report its statistical variables which are generated by grouping local features and re-encoding the instances. The candidate parties receive the statistical variables and score each feature based on Conditional Mutual Information (CMI). The feature scores are then collected by an active party to perform global ranking to produce the best feature subset $S_{k}$ (s.t., $|S_{k}|=m$ given $m$ as the number of local features to select) for each client $u_{k}$, which collectively represents the solution to the following optimization:
$\min_{S_{k}\subseteq\mathcal{X}_{k}}H(Y|\{S_{1},S_{2},\ldots\}).$ (17)
@@
where $H(Y|\{S_{1},S_{2},\ldots\})$ is conditional entropy of the label $Y$ given the union of locally selected feature subsets $\{S_{1},S_{2},\ldots\}$. This framework offers useful insights for communication-efficient exchange of knowledge for collaborative feature selection, but the protocol does not consider dishonest parties (clients) in the system. As a matter of fact, feature engineering techniques can be utilized to defend against data poisoning attacks strengthened by feature manipulation. This is also termed feature poisoning (FP) in the literature [101]. Experimental evidence has demonstrated that, despite being highly disguised, FP attacks can be effectively countered by the Random Deep Feature Selection (RDFS) [102] strategy.
@@
Beside feature selection, feature transformation represents another typical approach to data engineering. By mapping samples to a new space of features derived from raw data, feature transformation not only helps to reduce information redundancy but also refines the representation of knowledge that is difficult to learn. However, it is not trivial for data holders in a federated system to perform feature transformation collaboratively. A practical problem that stands out is the computation cost incurred by performance evaluation provided numerous combinations of transformed features. Moreover, privacy issues arise since transformation can involve multiple features coming from different domains, which necessitates Secure Multi-Party Computation (MPC). To address these issues, Fang et al. [103] designed FLFE, a federated feature engineering framework for efficient, privacy-preserving multi-party collaboration. The core techniques employed include a sketching scheme for quantizing local features, a DNN-based classifier for judging the value of features, and a mask-based feature exchange protocol for transformation. This brings advantages in communication efficiency and also preserves data privacy without homomorphic encryption that limits the diversity of transformations.
@@
TABLE IV: Summarizing the studies related to federated label and feature space correction in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
@@
[88, 89] Fed. label correction local intrinsic dimensionality (LID) and models $\bullet$ global pre-processing plus “clean” clients-only fine-tuning
@@
[84] Fed. label correction warm-up models and class-wise loss vectors $\bullet$ distance-aware aggregation to mitigate noisy knowledge
@@
[95] Fed. label correction class-wise confidence thresholds and models $\bullet$ consensus-based method to ensemble decentralized knowledge $\bullet$ reliance on a pre-trained model for label uncertainty estimation MNIST, USC-HAD, CIFAR-10, and Clothing1M
[97] Fed. label correction denoising rate decisions and models $\bullet$ payoff modeling for game-based interaction
@@
[98, 99] Fed. feature selection MI-based scores and feature subsets $\bullet$ low communication costs
@@
[100] Fed. feature selection statistical variables and feature scores $\bullet$ MI-based and model-independent
@@
[104] Fed. feature selection feature subsets $\bullet$ low communication cost
@@
[105] Fed. feature transform. expression of derived features $\bullet$ low communication cost and reduced search space $\bullet$ very high cost of feature subset evaluation OpenML586 and Airfoil
[106] Fed. feature selection probability vectors associated with features $\bullet$ probabilistic feature subset selection and Bayesian aggregation
@@
[103] Fed. feature transform. sketches of features $\bullet$ Quantile Data Sketch and mask-based feature exchange
@@
A major obstacle for adapting Automated Feature Engineering (AutoFE) methods (e.g., DIFER [107], OpenFE [108] and Cognito [109]) to the federated system is the prohibitive cost of communication. To this end, Overman and Klabjan [105] have made early effort and devised a federated framework for collaborative AutoFE. In the horizontal FL setting, clients independently carry out AutoFE and report the expression of derived features to the server. The server then randomly applies masks to the union of these features to generate feature subsets. An iterative halving strategy is used to find the best mask (e.g., optimal feature subset) after running multiple rounds of standard FL for performance evaluation. For vertical FL setting, feature values across the entire local dataset need to be shared after homomorphic encryption so as to fuse features from different domains. For hybrid settings, importance of common features across multiple domains can be assessed progressively with the training process.
Federated feature engineering can facilitate the refinery of knowledge in a broad range of scenarios. For example, Zhang et al. [104] proposed a federated feature selection framework for Internet of Things (IoT) applications where each device (as a client) has very limited computational resources. In this framework, local features $\{f^{(k)}_{1},f^{(k)}_{2},\ldots\}$ on each client $u_{k}$ are first filtered using a threshold defined based on Symmetric Uncertainty (SU) and Feature Average Relevance (FAR):
@@
where $d_{k}$ is the feature dimensionality for client $u_{k}$, $H(\cdot)$ is the entropy measure and $MI(\cdot,\cdot)$ denotes the mutual information between two features. Next, One-Class Support Vector Machine (OCSVM) and hierarchical clustering are employed on each client to identify the subset of most significant features. The server then calculates the intersection of these subsets and synchronizes with the clients. The solution was evaluated on four classification tasks in relevance to IoT applications. Experimental results demonstrated that competitive and even higher performance could be achieved with the federated feature subset, with improved training efficiency.
Cassará et al. [106] argued that a small subset of features are adequate for collaborative learning in interconnected Autonomous Driving Systems (ADSs). To this end, they developed a federated feature selection algorithm that combines on-vehicle feature filtering and edge-based feature subset aggregation. Locally, each participating vehicle (as a client) optimizes a feature selection probability vector such that the conditional entropy $H(y|S)$ (given feature subset $S$ and labels $y$) is minimized, which corresponds to an Associated Stochastic Problem (ASP). The edge server then aggregates these probability vectors into a consensus one and feeds it back to the vehicles. This interaction keeps going iteratively until the global vector converges. The Kolmogorov-Smirnov test with preset parameters are used for checking convergence. By evaluating this algorithm, the authors claimed that a 99% compression rate for network traffic was achieved since as few as 24 (out of 2166) features were selected for collaborative training on the MAV dataset, with negligible impact on model performance. This finding indicates that federated feature engineering can be as effective as model-based compression methods (e.g., quantization, pruning, etc.) [110] in terms of saving learning costs.

### v3

@@ V-D Summary
With rising concerns of data-centric attacks, dataset-level filtering can minimize the risk of involving poisonous data. We argue that fine-grained filtering at sample level can be employed jointly with participant screening to provide more reliable evidence for adversary detection. As of data re-balancing, existing approaches stand on an implicit assumption that uniform data distribution is ideal for the downstream task. However, this may not always hold. On the one hand, the local learner could under-fit on the “special knowledge” [tang2021data]. On the other hand, there is no sufficient theoretical evidence that supports imbalanced sampling, provided that the test data distribution is unknown. Besides, collaborative data refinement for regression tasks, e.g. drug response modeling, is hardly explored. Moreover, existing studies pay little attention to the subtle line between data imbalance and non-IIDness of data, the understanding of which can help us better redefine how data can refined locally and globally.
VI Collective Harmonization of Data
This section revisits branches of studies that employ collaborative mechanisms to make correction or transformation on data labels or features so that they can be utilized later in a harmonized fashion.
@@ VI-A Federated Data Correction
Mislabeled samples can have salient impact. This motivates researches on federated noisy label learning (FNLL) [jiang2024tackling, wu2023fednoro]. Considering the tie between a participant and its local domain, Xu et al. [xu2022fedcorr] proposed FedCorr, a multi-stage label denoising framework that streamlines a pre-processing stage and a fine-tuning stage prior to the traditional learning process, all executed in a federated fashion. In the pre-processing stage, each participant computes an estimate of the Local Intrinsic Dimensionality (LID) based on its local model’s predictions, which are collected by the central server for noisy dataset detection. As an useful measure that connects dimensionality and discriminability [houle2013dimensionality, karger2002finding], the LID for a vector $\boldsymbol{x}\sim\mathcal{D}$ is defined as:
@@
where $d_{x}$ is the distance between $\boldsymbol{x}$ and a randomly drawn vector $\boldsymbol{x}^{\prime}$ from $\mathcal{D}$ and $F_{d_{x}}(\cdot)$ represents the cumulative distribution function (CDF) of $d_{x}$. As models trained on noisy data tend to report larger LID scores, the server then uses this rule to spot these participants for label correction.
FedCorr [xu2022fedcorr] adopts a confidence-based rule to find and correct mislabeled data on noisy datasets. Specifically, the subset of samples (controlled by a proportion $\pi$) that yield the highest losses is first selected, following which the predictions of the global model $f_{G}$ are used for relabeling. Formally, the final subset of samples $D^{\prime}_{k}$ (for a noisy client $u_{k}$) to be relabeled is defined in (9):
$D^{\prime}_{k}=\big\{(\boldsymbol{x},y)|(\boldsymbol{x},y)\in\\ \underset{\hat{D}\subseteq D_{k}\atop|\hat{D}|=\pi|D_{k}|}{\arg\max}\sum_{(\boldsymbol{x},y)\in\hat{D}}\ell(f_{G}(\boldsymbol{x}),y)\text{ and }\max f_{G}(\boldsymbol{x})\geq\theta\big\},$ (9)
where $\ell$ is the loss function and $D_{k}$ is the local data from participant $u_{k}$. The labels of $D^{\prime}_{k}$ are then corrected to the predictions by the global model $f_{G}$.
There are two major defects in FedCorr [xu2022fedcorr]: i) incompatibility to general loss functions other than cross entropy, and ii) susceptibility to dishonest participants who tamper with their LID scores. Targeting at the second issue, Wang et al. [wang2024icde] built upon FedCorr and redesigned its federated pre-processing stage to achieve robust label correction. The key idea behind is to craft an efficient Zero-Knowledge Proof (ZKP) [campanelli2022zkp] protocol by which the server (as the verifier) can verify the computation integrity of the LID scores reported by the participants (as the provers) in a batched manner. The proofs are generated both before and after label correction. Only calculation in the form of arithmetic circuits is verifiable by the protocol, which limits its potential application to other data correction schemes. Similarly, the FedELC framework by Jiang et al. [jiang2024tackling] detects participants that carry noisy data and corrects them in two separate stages. The first stage is almost the same as in [xu2022fedcorr, wu2023fednoro], whilst the second stage creates learnable soft labels for data correction optimized towards the minimization of a linear combination of three different entropy losses. FedELC was evaluated on a rich set of noise data settings including the CIFAR-N [wei2021cifar-n] dataset from Amazon Mechanical Turk which features a high noise rate of human-annotated labels.
Noisy label detection and correction can also be executed directly at sample level. Based on label uncertainty estimation [northcutt2021confident], Zeng et al. [zeng2022clc] started from utilizing a well-trained global model’s predictions to estimate the probability $P_{k,i}$ that samples of a class $i\in\mathcal{C}$ are correctly labeled in a local dataset $D_{k}$, given in (10):
@@
where $D_{k,y=i}$ is the data subset on participant $u_{k}$ labeled as class $i$, $\mathcal{C}\subset\mathbb{W}$ is the label space, and $\theta_{G}$ denotes the global model’s parameter. Next, the global estimate of the probability of correctly-labeled class $i$ can be obtained using weighted average consensus [zeng2022clc] over the population $\mathcal{K}$:
@@
which resembles the model aggregation rule in the canonical FedAvg [mcmahanfl].
A practical concern of federated data correction is that the participants can be self-interested, which means that they may not stick to the protocol if their gain of participation cannot cover the cost of data correction. In light of this issue, Yan et al. [yan2024incent] first modeled the payoff of participants as a function of the model’s accuracy, payment from the task publisher, and the data correction cost. The authors employed an iterative denoising strategy that corrects a fraction of $\alpha$ noise samples per round. To reduce the fraction of noisy data from $\epsilon$ to $\epsilon^{\prime}$ given a denoising cost per sample $c$ and data size $|D_{k}|$, the total cost of data correction for client $u_{k}$ is:
@@
Yan et al. [yan2024incent] considered two different situations of federated denoising. In the first case, participants are cooperative and work towards to the optimal social welfare, i.e., the maximization of the sum of client payoffs. In the second case, every participant is assumed self-interested and only motivated to maximize its own payoff, which corresponds to a non-cooperative game.
@@ VI-B Federated Feature Engineering
In TAF, by federated feature engineering we refer to a new thread of research with particular focus on aligning, selecting and transforming the feature space (Fig. 6). This process can either be independent of the downstream task or somehow coupled with a specific protocol, motivated to address three challenges:
  • Disordered or misaligned attributes: in cross-institutional cooperation it is common to deal with different views of data where the local datasets are subject to different permutations of attributes [thakur2024knowledge].
@@
These challenges have catalyzed cutting-edge studies on new feature engineering paradigms. In this section we discuss the advance in federated feature engineering, with particular focuses on collaborative feature selection and feature transformation.
Figure 6: A schematic illustration of federated feature engineering.
Fed-FiS [banerjee2021fedfis] represents an early exploration of collaborative feature selection in the federation based on information theory. The study follows a training-free, task-agnostic approach to assessing the importance of features. Locally, a participant $u_{k}$ independently evaluates the feature-feature mutual information (FFMI) and the feature-class mutual information (FCMI) over the local feature and label set $\{f^{(k)}_{1},f^{(k)}_{2},\ldots\}\cup C^{(k)}$ by Eqs. (13) and (14):
@@
Based on these mutual information-based measures, each participant performs clustering to select a subset of features which are then uploaded to the central server. In the second stage, the server uses FFMI and FCMI again to score each feature, which in turn helps the participants identify the best local feature subset. Banerjee et al. [banerjee2024cost] further extended this framework using an improved ranking scheme based on Pareto optimization where each feature is treated as a solution characterized by FFMI and FCMI. These solutions are then ranked based on how many others dominate them and that they dominate. Rankings are reported back to the clients.
@@
Fu et al. [fu2023feast] studied the feature selection problem in the Vertical Federated Learning (VFL) setting where the set of entities are associated with consistent identifiers but represented by different features. The proposed framework, FEAST, involves three types of participants for collaborative feature selection: a selected party, an active party, and candidate parties. In the pre-processing stage, discretization is first applied to each feature, combined with stratified sampling to reduce computation and communication costs. This is followed by iterative operations and interactions between the parties. Specifically, a party is selected to report its statistical variables which are generated by grouping local features and re-encoding the instances. The candidate parties receive the statistical variables and score each feature based on Conditional Mutual Information (CMI) [yang1999feature]. The feature scores are then collected by an active party to perform global ranking to produce the best feature subset $S_{k}$ (s.t., $|S_{k}|=m$ given $m$ as the number of local features to select) for each participant $u_{k}$, which collectively represents the solution to the following optimization:
$\min_{S_{k}\subseteq\mathcal{X}_{k}}H(Y|\{S_{1},S_{2},\ldots\}),$ (15)
@@
where $H(Y|\{S_{1},S_{2},\ldots\})$ is the conditional entropy of the label $Y$ given the union of locally selected feature subsets $\{S_{1},S_{2},\ldots\}$. This framework offers useful insights for communication-efficient exchange of knowledge for collaborative feature selection, but the protocol does not consider dishonest participants in the system. As a matter of fact, feature engineering techniques can be utilized to defend against data poisoning attacks strengthened by feature manipulation. This is also termed feature poisoning (FP) in the literature [Nowroozi2024rdfs]. Experimental evidence has demonstrated that, despite being highly disguised, FP attacks can be effectively countered by the Random Deep Feature Selection (RDFS) [Nowroozi2025fp] strategy.
@@
Beside feature selection, feature transformation represents another typical approach to data engineering. By mapping samples to a new space of features derived from raw data, feature transformation not only helps to reduce information redundancy but also refines the representation of knowledge that is difficult to learn. However, it is not trivial for data holders in a federated system to perform feature transformation collaboratively. A practical problem that stands out is the computation cost incurred by performance evaluation given numerous combinations of transformed features. Moreover, privacy issues arise since transformation can involve multiple features coming from different domains, which necessitates Secure Multi-Party Computation (MPC). To address these issues, Fang et al. [fang2020flfe] designed FLFE, a federated feature engineering framework for efficient, privacy-preserving multi-party collaboration. The core techniques employed include a sketching scheme for quantizing local features, a DNN-based classifier for judging the value of features, and a mask-based feature exchange protocol for transformation. This brings advantages in communication efficiency and also preserves data privacy without homomorphic encryption that limits the diversity of transformations.
@@
TABLE IV: Summarizing the studies related to collective harmonization of data in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
@@
[xu2022fedcorr, wang2024icde] Federated label correction local intrinsic dimensionality (LID) and models $\bullet$ global pre-processing plus “clean” clients-only fine-tuning
@@
[jiang2024tackling] Federated label correction warm-up models and class-wise loss vectors $\bullet$ distance-aware aggregation to mitigate noisy knowledge
@@
[zeng2022clc] Federated label correction class-wise confidence thresholds and models $\bullet$ consensus-based method to ensemble decentralized knowledge $\bullet$ reliance on a pre-trained model for label uncertainty estimation MNIST, USC-HAD, CIFAR-10, and Clothing1M
[yan2024incent] Federated label correction denoising rate decisions and models $\bullet$ payoff modeling for game-based interaction
@@
[banerjee2021fedfis, banerjee2024cost] Federated feature selection MI-based scores and feature subsets $\bullet$ low communication costs
@@
[fu2023feast] Federated feature selection statistical variables and feature scores $\bullet$ MI-based and model-independent
@@
[zhang2023fshfl] Federated feature selection feature subsets $\bullet$ low communication cost
@@
[cassara2022tvt] Federated feature selection probability vectors associated with features $\bullet$ probabilistic feature subset selection and Bayesian aggregation
@@
[overman2024federated] Federated feature transform expression of derived features $\bullet$ low communication cost and reduced search space $\bullet$ very high cost of feature subset evaluation OpenML586 and Airfoil
[fang2020flfe] Federated feature transform sketches of features $\bullet$ Quantile Data Sketch and mask-based feature exchange
@@
A major obstacle for adapting Automated Feature Engineering (AutoFE) methods (e.g., OpenFE [zhang2023openfe]) to the federated system is the prohibitive cost of communication. To this end, Overman and Klabjan [overman2024federated] have made early effort and devised a federated framework for collaborative AutoFE. In their framework, participants independently carry out AutoFE and report the expression of derived features to the server. The server then randomly applies masks to the union of these features to generate feature subsets. An iterative halving strategy is used to find the best mask (e.g., optimal feature subset) after running multiple rounds of learning for performance evaluation. For participants with heterogeneous feature space, feature values across the entire local datasets need to be shared after homomorphic encryption so as to fuse features from different domains. For hybrid settings, importance of common features across multiple domains can be assessed progressively with the process.
Federated feature engineering can facilitate the refinement of knowledge in a broad range of scenarios. For example, Zhang et al. [zhang2023fshfl] proposed a federated feature selection framework for Internet of Things (IoT) applications where each device (as a participant) has very limited computational resources. In this framework, local features $\{f^{(k)}_{1},f^{(k)}_{2}$$\ldots\}$ on each client $u_{k}$ are first filtered using a threshold defined based on Symmetric Uncertainty (SU) and Feature Average Relevance (FAR):
@@
where $d_{k}$ is the feature dimensionality for participant $u_{k}$, $H(\cdot)$ is the entropy measure and $MI(\cdot,\cdot)$ denotes the mutual information between two features. Next, One-Class Support Vector Machine (OCSVM) and hierarchical clustering are employed on each participant to identify the subset of most significant features. The server then calculates the intersection of these subsets and synchronizes with the participants. The solution was evaluated on four classification tasks in relevance to IoT applications. Experimental results demonstrated that competitive and even higher performance could be achieved with the federated feature subset, with improved efficiency.
Cassará et al. [cassara2022tvt] argued that a small subset of features are adequate for an interconnected Autonomous Driving Systems (ADSs). To this end, they developed a federated feature selection algorithm that combines on-vehicle feature filtering and edge-based feature subset aggregation. Locally, each participating vehicle optimizes a feature selection probability vector such that the conditional entropy $H(y|S)$ (given feature subset $S$ and labels $y$) is minimized, which corresponds to an Associated Stochastic Problem (ASP). The edge server then aggregates these probability vectors into a consensus one and feeds it back to the vehicles. This interaction keeps going iteratively until the global vector converges. The Kolmogorov-Smirnov test with preset parameters are used for checking convergence. By evaluating this algorithm, the authors claimed that a 99% compression rate for network traffic was achieved since as few as 24 (out of 2166) features were selected for collaborative training on the MAV dataset, with negligible impact on downstream tasks. This finding indicates that federated feature engineering can be as effective as model-based compression methods (e.g., quantization, pruning, etc.) [gupta2022compression] in terms of cost saving.